# Useful stochastic bounds in time-varying queues with service and patience times having general joint distribution

### v1

@@ 2.3 Proof of Theorem 1
The next step is to derive an upper bound for the limit that appears in the right-hand side of (2.3). To this end, we will need the following lemma.
@@
Now, recall that: (1) The service discipline is LCFS-PR. (2) The arrival process is homogeneous Poisson process with rate $\lambda_{h}$. Then, using the strong Markov property of the workload process with the memoryless property of the exponential distribution, we deduce that the lower bound in the right-hand side of (39) is distributed like $A^{(k)}_{\lambda_{h}}(g;x,\Psi)$, from which Lemma 1 follows.
@@
$\lim_{k\uparrow\infty}\mathbb{E}f\left[\int_{0}^{\tau_{\lambda_{h}}^{(k)}}g\circ W_{\lambda_{h}}^{(k)}(t){\rm d}t\right]$ $=\lim_{k\uparrow\infty}\mathbb{E}A_{k}=\mathbb{E}\sup_{k\in\mathbb{Z}_{\geq 0}}A_{k}$ (40)
@@
$\mathbb{E}f\left[\int_{0}^{\tau_{\lambda}^{(\infty)}}g_{n}\circ W_{\lambda}^{(\infty)}(t){\rm d}t\right]\leq\mathbb{E}f\left[\int_{0}^{\tau_{\lambda_{h}}^{(\infty)}}g_{n}\circ W_{\lambda_{h}}^{(\infty)}(t){\rm d}t\right]$ (42)
@@
$\mathbb{P}\left\{A_{\lambda}(g)\geq s\right\}\leq\mathbb{P}\left\{A_{\lambda_{h}}(g)\geq s\right\}$ (44)
@@ 3 Proof of Theorem 3
$\iota_{\lambda}\equiv\inf\left\{i\in\mathbb{N};|I_{\lambda}^{i}|>\kappa\right\}$ (45)
@@
$\zeta_{\lambda}\equiv\min\left\{\kappa n\in I_{\lambda}^{\iota_{\lambda}};W_{\lambda}(\kappa n)=0\ ,\ n\in\mathbb{N}\right\}\,.$ (46)
@@ 3.1 Proof of Proposition 2
$I^{i}_{\lambda_{h}}\equiv D_{\lambda_{h}}\left(\inf I_{\lambda}^{i}\right)-\inf I_{\lambda}^{i}\,.$ (51)
@@
$Q_{i}(s,t)\equiv Q^{-1}_{\lambda(t+\cdot))}\left(U_{i};s,\Psi\right)\ \ \left(\text{resp., }R_{i}(s)\equiv R^{-1}_{\lambda_{h}}\left(U_{i};s,\Psi\right)\right).$ (53)
@@
$\overline{\mathcal{A}}_{\lambda}(g)\stackrel{{\scriptstyle\text{d}}}{{=}}\iota_{\lambda}\kappa g(0)+\sum_{i=1}^{\iota_{\lambda}-1}Q\left(\sigma_{\lambda}^{i},\upsilon_{\lambda}^{i}\right)$ (55)
@@ 4 Proofs of Theorems 4 and 5
$A_{\lambda_{h}}(g;x,\Psi)\leq_{\text{st}}A_{\lambda_{h}}(g;x,\Psi_{\infty}).$ (57)
@@ 4.2 Proof of Theorem 4 and Theorem 5 (continuation)
$\mathbb{P}\left\{\mathcal{A}_{\lambda}(g)>u\right\}\leq\mathbb{P}\left\{\sum_{i=1}^{\iota_{\lambda_{h}}}\left[\widetilde{A}_{\lambda_{h}}^{i}+g(0)\max\{\kappa,1\}\right]>u\right\}\ \ ,\ \ \forall u>0\,,$ (58)
@@
$\lim_{u\to\infty}\frac{\mathbb{P}\left\{\sum_{i=1}^{\iota_{\lambda_{h}}}\left[\widetilde{A}_{\lambda_{h}}^{i}+g(0)\max\{\kappa,1\}\right]>u\right\}}{\mathbb{P}\left\{\widetilde{A}_{\lambda_{h}}^{1}>u-g(0)\max\{\kappa,1\}\right\}}=\frac{1-e^{-\kappa\lambda_{h}}}{e^{-\kappa\lambda_{h}}}\,.$ (59)
@@ 5 Proof of Theorem 6
$\mu_{\lambda_{h}}\equiv\int_{0}^{\tau_{\lambda_{h}}\left(S_{0}\right)}{\rm d}N^{*}_{\lambda_{h}}(t)$ (60)
@@
$\int_{0}^{\infty}y{\rm d}G_{\varepsilon}(y)=\varepsilon+\int_{0}^{\infty}y{\rm d}G(y)<\infty,$ (62)
@@
$\tau_{\lambda_{h},\varepsilon}\equiv\inf\left\{t\geq 0;W_{\lambda_{h},\epsilon}(t)=0\right\}\,.$ (63)
@@
$\int_{0}^{\tau_{\lambda_{h}}(S_{0})}g\circ W_{\lambda_{h}}(t;S_{0}){\rm d}N^{*}_{\lambda_{h}}(t)\leq_{\text{st}}\int_{0}^{\tau_{\lambda_{h},\varepsilon}(S_{0}+\varepsilon)}g\circ W_{\lambda_{h},\varepsilon}(t;S_{0}+\varepsilon){\rm d}N^{*}_{\lambda_{h},\varepsilon}(t)\,.$ (64)
@@
$\mu_{\lambda_{h},\varepsilon}\equiv\int_{0}^{\tau_{\lambda_{h},\varepsilon}}{\rm d}N^{*}_{\lambda_{h},\varepsilon}(t)$ (65)
@@
$0\leq\varepsilon\mu_{\lambda_{h},\varepsilon}\leq\tau_{\lambda_{h},\varepsilon}\,.$ (66)
@@ 5.1 Proof of Lemma 3
$\overline{E}_{i,\varepsilon}\equiv\inf\left\{t\geq E_{i-1,\varepsilon};W_{\lambda_{h},\varepsilon}(t)-W_{\lambda_{h},\varepsilon}(t-)>0\right\}\,,$ (67)
@@
$\left\{W_{\lambda_{h}}(t;S_{0});0\leq t\leq\tau_{\lambda_{h}}(S_{0})\right\}\stackrel{{\scriptstyle\text{d}}}{{=}}\left\{W_{\lambda_{h},\epsilon}(t);t\in\mathcal{T}_{\lambda_{h},\varepsilon}\right\}\,,$ (68)
@@
$\mathcal{T}_{\lambda_{h},\varepsilon}\equiv\left[0,\tau_{\lambda_{h}}(S_{0})\right]\setminus\bigcup_{i=0}^{\infty}\left(E_{i,\varepsilon},\overline{E}_{i,\varepsilon}\right]\,.$ (69)
@@
$\int_{0}^{\tau_{\lambda_{h},\varepsilon}}g\circ W_{\lambda_{h},\varepsilon}(t){\rm d}N^{*}_{\lambda_{h}}(t)$ $\geq\int_{\mathcal{T}_{\lambda_{h},\varepsilon}}g\circ W_{\lambda_{h},\varepsilon}(t){\rm d}N^{*}_{\lambda_{h}}(t)$ (70)
@@ 5.2 The proof of Lemma 4
$\mathbb{E}W_{1}(x)-x+1\leq\lambda_{h}\int_{0}^{\infty}y{\rm d}G(y)\left[1-H(x-1)\right]\,.$ (71)
@@
$W_{1}(x)-x+1=\sum_{i=1}^{\eta_{\lambda_{h}}^{*}\left([0,1];x\right)}S_{i}\,.$ (72)
@@
$\eta_{\lambda_{h}}^{*}\left([0,1];x\right)\leq\eta_{\lambda_{h}}^{**}\left([0,1];x\right)$ (73)
@@
$\mathbb{E}\sum_{i=1}^{\eta_{\lambda_{h}}^{*}\left([0,1];x\right)}S_{i}\leq\mathbb{E}\eta_{\lambda_{h}}^{**}\left([0,1];x\right)\int_{0}^{\infty}y{\rm d}G(y).$ (75)
@@
$\mathbb{E}\eta_{\lambda_{h}}^{**}\left([0,1];x\right)=\lambda_{h}\left[1-H(x-1)\right]\,,$ (76)
@@
$\lambda_{h}\left[1-H(k-1)\right]\int_{0}^{\infty}y{\rm d}G(y)<1-\delta\,.$ (77)
@@
$\psi(x)\equiv 1+\frac{x}{\delta}\ \ ,\ \ \forall x\geq 0\,.$ (78)
@@
$\left|\psi\left(W_{1}(x)\right)-\psi\left(x\right)\right|\leq\frac{|L_{1}(x)|+|T_{1}(x)|}{\delta}=\frac{L_{1}(x)+1}{\delta}\,.$ (81)
@@
$\sup_{x\in[0,\infty)}\mathbb{E}\left|\psi\left(W_{1}(x)\right)-\psi\left(x\right)\right|^{m}<\infty\,.$ (82)
@@
$\tau_{\mathcal{C}}(x)\equiv\tau^{(1)}_{\mathcal{C}}(x)\equiv\inf\left\{n\in\mathbb{Z}_{\geq 0};W_{n}(x)\in\mathcal{C}\right\}\,.$ (83)
According to Proposition 7.12(i) from the book by Benaïm and Hurth [9], since (79), (80) and (82) hold, there exists $\alpha>0$ (uniform in $x$) such that for every $x\in\mathcal{C}$:
@@
$\tau^{(n+1)}_{\mathcal{C}}(x)\equiv\inf\left\{n>\tau^{(n)}_{\mathcal{C}}(x);W_{n}(x)\in\mathcal{C}\right\}\,.$ (85)
@@
$\mathbb{P}\left\{\exists 1\leq i\leq k;Y_{i}=0\right\}\geq\mathbb{P}\left\{\text{No arrivals during}\ \left[0,k\right]\right\}\geq e^{-\lambda_{h}k}\,.$ (86)
@@ 6 Discussion and future research
$A_{\lambda_{1}}(g;x,\Psi)\leq_{\text{st}}A_{\lambda_{2}}(g;x,\Psi),\ \forall x>0,\ g\in\mathcal{G},\ \Psi\in\mathcal{L}\,.$ (87)
@@
Our Theorem 1 includes an affirmative answer when $\lambda_{2}(\cdot)$ is a constant function. Actually, by very minor modifications of the proof of Theorem 1, one may get an affirmative answer also when $\lambda_{1}(\cdot)$ is a constant function. However, it is still an open problem whether this is true when both $\lambda_{1}(\cdot)$ and $\lambda_{2}(\cdot)$ are time-varying. Obviously, in the current proof of Theorem 1, the memoryless property of the exponential distribution has a significant role in the coupling construction. Therefore, we do not see how the current proof may straightforwardly generalize to the more general case.
Tail distribution of $\mathcal{A}_{\lambda}(g)$ and $\mathcal{A}^{*}_{\lambda}(g)$

### v2

@@ 2.2 Proof of Theorem 1
The next step is to derive an upper bound for the limit that appears on the right-hand side of (2.2). To this end, we will need the following lemma.
@@
Now, recall that: (1) The service discipline is LCFS-PR. (2) The arrival process is homogeneous Poisson process with rate $\lambda_{h}$. Then, using the strong Markov property of the workload process with the memoryless property of the exponential distribution, we deduce that the lower bound on the right-hand side of (14) is distributed like $A^{(k)}_{\lambda_{h}}(g;x,\Psi)$, from which Lemma 1 follows.
@@
$\lim_{k\uparrow\infty}\mathbb{E}f\left[\int_{0}^{\tau_{\lambda_{h}}^{(k)}}g\circ W_{\lambda_{h}}^{(k)}(t){\rm d}t\right]$ $=\lim_{k\uparrow\infty}\mathbb{E}A_{k}=\mathbb{E}\sup_{k\in\mathbb{Z}_{\geq 0}}A_{k}$
@@
$\mathbb{E}f\left[\int_{0}^{\tau_{\lambda}^{(\infty)}}g_{n}\circ W_{\lambda}^{(\infty)}(t){\rm d}t\right]\leq\mathbb{E}f\left[\int_{0}^{\tau_{\lambda_{h}}^{(\infty)}}g_{n}\circ W_{\lambda_{h}}^{(\infty)}(t){\rm d}t\right]$
@@
$\mathbb{P}\left\{A_{\lambda}(g)\geq s\right\}\leq\mathbb{P}\left\{A_{\lambda_{h}}(g)\geq s\right\}$
@@ 3 Proof of Theorem 3
$\iota_{\lambda}\equiv\inf\left\{i\in\mathbb{N};|I_{\lambda}^{i}|>\kappa\right\}$
@@
$\zeta_{\lambda}\equiv\min\left\{\kappa n\in I_{\lambda}^{\iota_{\lambda}};W_{\lambda}(\kappa n)=0\ ,\ n\in\mathbb{N}\right\}\,.$
@@ 3.1 Proof of Proposition 2
$I^{i}_{\lambda_{h}}\equiv D_{\lambda_{h}}\left(\inf I_{\lambda}^{i}\right)-\inf I_{\lambda}^{i}\,.$
@@
$Q_{i}(s,t)\equiv Q^{-1}_{\lambda(t+\cdot))}\left(U_{i};s,\Psi\right)\ \ \left(\text{resp., }R_{i}(s)\equiv R^{-1}_{\lambda_{h}}\left(U_{i};s,\Psi\right)\right).$
@@
$\overline{\mathcal{A}}_{\lambda}(g)\stackrel{{\scriptstyle\text{d}}}{{=}}\iota_{\lambda}\kappa g(0)+\sum_{i=1}^{\iota_{\lambda}-1}Q\left(\sigma_{\lambda}^{i},\upsilon_{\lambda}^{i}\right)$
@@ 4 Proofs of Theorems 4 and 5
$A_{\lambda_{h}}(g;x,\Psi)\leq_{\text{st}}A_{\lambda_{h}}(g;x,\Psi_{\infty}).$
@@ 4.2 Proof of Theorem 4 and Theorem 5 (continuation)
$\mathbb{P}\left\{\mathcal{A}_{\lambda}(g)>u\right\}\leq\mathbb{P}\left\{\sum_{i=1}^{\iota_{\lambda_{h}}}\left[\widetilde{A}_{\lambda_{h}}^{i}+g(0)\max\{\kappa,1\}\right]>u\right\}\ \ ,\ \ \forall u>0\,,$
@@
$\lim_{u\to\infty}\frac{\mathbb{P}\left\{\sum_{i=1}^{\iota_{\lambda_{h}}}\left[\widetilde{A}_{\lambda_{h}}^{i}+g(0)\max\{\kappa,1\}\right]>u\right\}}{\mathbb{P}\left\{\widetilde{A}_{\lambda_{h}}^{1}>u-g(0)\max\{\kappa,1\}\right\}}=\frac{1-e^{-\kappa\lambda_{h}}}{e^{-\kappa\lambda_{h}}}\,.$
@@ 5 Proof of Theorem 6
$\mu_{\lambda_{h}}\equiv\int_{0}^{\tau_{\lambda_{h}}\left(S_{0}\right)}{\rm d}N^{*}_{\lambda_{h}}(t)$
@@
$\int_{0}^{\infty}y{\rm d}G_{\varepsilon}(y)=\varepsilon+\int_{0}^{\infty}y{\rm d}G(y)<\infty,$
@@
$\tau_{\lambda_{h},\varepsilon}\equiv\inf\left\{t\geq 0;W_{\lambda_{h},\epsilon}(t)=0\right\}\,.$
@@
$\int_{0}^{\tau_{\lambda_{h}}(S_{0})}g\circ W_{\lambda_{h}}(t;S_{0}){\rm d}N^{*}_{\lambda_{h}}(t)\leq_{\text{st}}\int_{0}^{\tau_{\lambda_{h},\varepsilon}(S_{0}+\varepsilon)}g\circ W_{\lambda_{h},\varepsilon}(t;S_{0}+\varepsilon){\rm d}N^{*}_{\lambda_{h},\varepsilon}(t)\,.$
@@
$\mu_{\lambda_{h},\varepsilon}\equiv\int_{0}^{\tau_{\lambda_{h},\varepsilon}}{\rm d}N^{*}_{\lambda_{h},\varepsilon}(t)$
@@
$0\leq\varepsilon\mu_{\lambda_{h},\varepsilon}\leq\tau_{\lambda_{h},\varepsilon}\,.$
@@ 5.1 Proof of Lemma 3
$\overline{E}_{i,\varepsilon}\equiv\inf\left\{t\geq E_{i-1,\varepsilon};W_{\lambda_{h},\varepsilon}(t)-W_{\lambda_{h},\varepsilon}(t-)>0\right\}\,,$
@@
$\left\{W_{\lambda_{h}}(t;S_{0});0\leq t\leq\tau_{\lambda_{h}}(S_{0})\right\}\stackrel{{\scriptstyle\text{d}}}{{=}}\left\{W_{\lambda_{h},\epsilon}(t);t\in\mathcal{T}_{\lambda_{h},\varepsilon}\right\}\,,$
@@
$\mathcal{T}_{\lambda_{h},\varepsilon}\equiv\left[0,\tau_{\lambda_{h}}(S_{0})\right]\setminus\bigcup_{i=0}^{\infty}\left(E_{i,\varepsilon},\overline{E}_{i,\varepsilon}\right]\,.$
@@
$\int_{0}^{\tau_{\lambda_{h},\varepsilon}}g\circ W_{\lambda_{h},\varepsilon}(t){\rm d}N^{*}_{\lambda_{h}}(t)$ $\geq\int_{\mathcal{T}_{\lambda_{h},\varepsilon}}g\circ W_{\lambda_{h},\varepsilon}(t){\rm d}N^{*}_{\lambda_{h}}(t)$
@@ 5.2 The proof of Lemma 4
$\mathbb{E}W_{1}(x)-x+1\leq\lambda_{h}\int_{0}^{\infty}y{\rm d}G(y)\left[1-H(x-1)\right]\,.$
@@
$W_{1}(x)-x+1=\sum_{i=1}^{\eta_{\lambda_{h}}^{*}\left([0,1];x\right)}S_{i}\,.$
@@
$\eta_{\lambda_{h}}^{*}\left([0,1];x\right)\leq\eta_{\lambda_{h}}^{**}\left([0,1];x\right)$
@@
$\mathbb{E}\sum_{i=1}^{\eta_{\lambda_{h}}^{*}\left([0,1];x\right)}S_{i}\leq\mathbb{E}\eta_{\lambda_{h}}^{**}\left([0,1];x\right)\int_{0}^{\infty}y{\rm d}G(y).$
@@
$\mathbb{E}\eta_{\lambda_{h}}^{**}\left([0,1];x\right)=\lambda_{h}\left[1-H(x-1)\right]\,,$
@@
$\lambda_{h}\left[1-H(k-1)\right]\int_{0}^{\infty}y{\rm d}G(y)<1-\delta\,.$
@@
$\psi(x)\equiv 1+\frac{x}{\delta}\ \ ,\ \ \forall x\geq 0\,.$
@@
$\left|\psi\left(W_{1}(x)\right)-\psi\left(x\right)\right|\leq\frac{|L_{1}(x)|+|T_{1}(x)|}{\delta}=\frac{L_{1}(x)+1}{\delta}\,.$
@@
$\sup_{x\in[0,\infty)}\mathbb{E}\left|\psi\left(W_{1}(x)\right)-\psi\left(x\right)\right|^{m}<\infty\,.$
@@
$\tau_{\mathcal{C}}(x)\equiv\tau^{(1)}_{\mathcal{C}}(x)\equiv\inf\left\{n\in\mathbb{Z}_{\geq 0};W_{n}(x)\in\mathcal{C}\right\}\,.$
According to Proposition 7.12(i) from the book by Benaïm and Hurth [9], since (23), (24) and (5.2) hold, there exists $\alpha>0$ (uniform in $x$) such that for every $x\in\mathcal{C}$:
@@
$\tau^{(n+1)}_{\mathcal{C}}(x)\equiv\inf\left\{n>\tau^{(n)}_{\mathcal{C}}(x);W_{n}(x)\in\mathcal{C}\right\}\,.$
@@
$\mathbb{P}\left\{\exists 1\leq i\leq k;Y_{i}=0\right\}\geq\mathbb{P}\left\{\text{No arrivals during}\ \left[0,k\right]\right\}\geq e^{-\lambda_{h}k}\,.$
@@ 6.1 Stochastic monotonicity:
$A_{\lambda_{1}}(g;x,\Psi)\leq_{\text{st}}A_{\lambda_{2}}(g;x,\Psi),\ \forall x>0,\ g\in\mathcal{G},\ \Psi\in\mathcal{L}\,.$
@@
Our Theorem 1 includes an affirmative answer when $\lambda_{2}(\cdot)$ is a constant function. Actually, by very minor modifications of the proof of Theorem 1, one may get an affirmative answer also when $\lambda_{1}(\cdot)$ is a constant function. However, it is still an open problem whether this is true when both $\lambda_{1}(\cdot)$ and $\lambda_{2}(\cdot)$ are time-varying.
In the current proof of Proposition 1, the memoryless property of the exponential distribution plays an important role in the coupling construction. Specifically, we need to couple two $\text{M}_{t}/\text{G}(\Psi)/1/k+\text{H}(\Psi)$ queues having, respectively, the arrival rates $\lambda_{1}(\cdot)$ and $\lambda_{2}(\cdot)$. Assume that $\lambda_{1}(t)\leq\lambda_{2}(t)=\lambda_{h}$ for every $t\geq 0$ and in both queues the total service time of the initial customer is $x$. This allowed us to apply the thinning argument (Theorem 1 from the work by Lewis and Shedler [34]) only on the service periods of the initial customer and argue that during all other times the arrivals are dictated by an external Poisson process with rate $\lambda_{h}$ independent of the arrival process during the service times of the initial customer. Due to the memoryless property of the exponential distribution (or equivalently, due to the strong Markov property of the homogeneous Poisson process), the resulting arrival process is still a Poisson process with rate $\lambda_{2}\equiv\lambda_{h}$. When $\lambda_{2}$ varies, we do not see how to adjust this construction of the arrival process to produce a Poisson arrival process with rate $\lambda_{2}(\cdot)$, i.e., the current proof may not be straightforwardly extended to the more general case.
6.2 Stochastic $\lambda(\cdot)$
It is intriguing to think whether Theorems 1 and 2 are valid when $\lambda(\cdot)$ is a non-negative stochastic process such that $\lambda(t)\leq\lambda_{h}$ for every $t\geq 0$, $\mathbb{P}$-a.s. If $\lambda$ is independent of the sequence $(S_{1},Y_{1}),(S_{2},Y_{2}),\ldots$, then Theorems 1 and 2 are easily extended by applying the same reasoning that leads to (6) (namely, conditioning and un-conditioning on $\lambda(\cdot)$). A class of models belonging to this setup is based on a continuous-time Markov chain $J(\cdot)$ with countable state-space $\mathcal{J}$ independent of $(S_{1},Y_{1}),(S_{2},Y_{2}),\ldots$ such that $\lambda(t)\equiv\lambda_{J(t)}(t)$ where $\lambda_{j}(\cdot),j\in\mathcal{J}$ are all deterministic rate functions bounded from above by $\lambda_{h}$. Note that under the additional assumption that $\lambda_{j}(\cdot),j\in\mathcal{J}$ are all constant functions (not necessarily identical ones), then the arrival process will become the so called Markov modulated Poisson process (see, e.g., the classic work by Neuts [38]).
It is also logical to think about state-dependent $\lambda(\cdot)$, i.e., when $\lambda(t)$ is modulated by the state of the queue. Note that conditioning and un-conditioning on $\lambda(\cdot)$ is not useful anymore and other solution methods should be applied, e.g., one may try to generalize the current proof.
For example, let $L(t)$ be the queue length process and assume that $\lambda(t)\equiv\lambda_{L(t)}(t)\leq\lambda_{h}$ for every $t\geq 0$ (with $\lambda_{\ell}(\cdot)$ a deterministic rate function for every $\ell\in\mathbb{Z}_{\geq 0}$). For this setup, one may verify that Proposition 1 remains valid, but then, when taking the size of the waiting room $k$ to infinity, a problem arises as the workload process in such a state-dependent queue is sensitive to the discipline of the service.
Another kind of potential state-dependence occurs under the assumption that $\lambda(\cdot)$ depends on the existing workload, i.e., if $T$ is distributed like the time until the next arrival (starting from initial workload $x$), then
$\mathbb{P}\left\{T>t\right\}=\exp\left\{-\int_{0}^{t}\lambda\left[t,W_{\lambda}(t;x)\right]{\rm d}t\right\}\ \ ,\ \ \forall t\in(0,\infty).$
For more details, see, e.g., the paper by Bekker et al. [8]. In this case, the workload process is invariant to the change of service discipline. However, the proof of Proposition 1 requires a generalization of the thinning arguement (i.e., Theorem 1 in the work by Lewis and Shedler [34]) to state-dependent Poisson arrival processes. To the best of our knowledge, extensions in this direction (of Theorem 1 in the work by Lewis and Shedler [34]) do not appear in the existing literature and, respectively, we consider its derivation as an interesting open question for future research.
6.3 Bounds on $\mathbb{E}g\left[W_{\lambda}(\infty)\right]$
Consider the periodic $\text{M}_{t}/\text{G}(\Psi)/1+\text{H}(\Psi)$ queue discussed in Section 1.4 with the notations presented in the proof of Theorem 3 and for simplicity assume that $g\in\mathcal{G}$ is such that $g(0)=0$. The regenerative property of the workload process yields, that when the queue is stable, we have the following identity of the steady-state workload:
$\mathbb{E}g\left[W_{\lambda}(\infty)\right]=\frac{\mathbb{E}\mathcal{A}_{\lambda}(g)}{\mathbb{E}\xi_{\lambda}}\,.$ (26)
Note that any customer who arrives at an empty queue will join the queue. Thus, under the assumption that the $\text{M}/\text{G}(\Psi)/1+\text{H}(\Psi)$ queue with an arrival rate $\lambda_{h}$ is also stable and $g(0)=0$, the regenerative property of the workload process yields:
$\mathbb{E}g\left[W_{\lambda_{h}}(\infty)\right]=\frac{\mathbb{E}A_{\lambda_{h}}(g)}{\frac{1}{\lambda_{h}}+\mathbb{E}\tau_{\lambda_{h}}}.$
It might be interesting to derive bounds for $\mathbb{E}g\left[W_{\lambda}(\infty)\right]$ in terms of $\mathbb{E}g\left[W_{\lambda_{h}}(\infty)\right]$. In practice, such a bound might be useful, e.g., when $\Psi$ has a product-form, because then the distribution of $W_{\lambda_{h}}(\infty)$ can be figuered out (at least for special cases) as described by Baccelli, Boyer, and Hebuterne [4].
To see, how such a bound may be constructed, recall the following result
$\mathcal{A}_{\lambda}(g)\leq_{\text{st}}\sum_{i=1}^{\iota_{\lambda_{h}}}A_{\lambda_{h}}^{i},$
which stems from Proposition 2. In addition, assume that $\lambda(t)\geq\lambda_{\ell}$ for every $t\geq 0$ for some $\lambda_{\ell}>0$. Since any customer who arrives to an empty queue, joins the queue, by construction (with the help of the strong Markov property) deduce that
$\xi_{\lambda}\geq_{\text{st}}\tau_{\lambda(\cdot+E)}\textbf{1}_{[0,\kappa]}(E),$
where:
  • $E$ is a random variable distributed like the arrival epoch of the first customer in the $\text{M}_{t}/\text{G}(\Psi)/1+\text{H}(\Psi)$ queue. Note that as $\lambda(t)\geq\lambda_{\ell}$ for every $t\geq 0$, we have:
    $\mathbb{P}\left\{E\leq\kappa\right\}\geq 1-e^{-\lambda_{\ell}\kappa}.$
  • For every $e\geq 0$, $\tau_{\lambda(\cdot+e)}$ is distributed as $\tau_{\widetilde{\lambda}}$ such that $\widetilde{\lambda}(t)=\lambda(t+e)$ for every $t\geq 0$.
  • The random variables belonging to the collection $\{\tau_{\lambda(\cdot+e)};e>0\}\cup\{E\}$ are mutually independent and also independent of the $\text{M}_{t}/\text{G}(\Psi)/1+\text{H}(\Psi)$ queue.
Now, since $\iota_{\lambda_{h}}\sim\text{Geo}(e^{-\lambda_{h}\kappa})$, by combining these results all together, deduce that
$\mathbb{E}g\left[W_{\lambda}(\infty)\right]$ $\leq e^{\lambda_{h}\kappa}\cdot\frac{\frac{1}{\lambda_{h}}+\mathbb{E}\tau_{\lambda_{h}}}{\mathbb{P}\left\{E\leq\kappa\right\}\mathbb{E}\left[\tau_{\lambda(\cdot-E^{\prime})}\big{|}E\leq\kappa\right]}\cdot\mathbb{E}g\left[W_{\lambda_{h}}(\infty)\right]$
$\leq\frac{e^{\lambda_{h}\kappa}}{1-e^{-\lambda_{\ell}\kappa}}\cdot\frac{\frac{1}{\lambda_{h}}+\mathbb{E}\tau_{\lambda_{h}}}{\mathbb{E}\left[\tau_{\lambda(\cdot-E^{\prime})}\big{|}E^{\prime}\leq\kappa\right]}\cdot\mathbb{E}g\left[W_{\lambda_{h}}(\infty)\right]$
$\leq\frac{e^{\lambda_{h}\kappa}}{1-e^{-\lambda_{\ell}\kappa}}\cdot\left(1+\rho_{h}^{-1}\right)\mathbb{E}g\left[W_{\lambda_{h}}(\infty)\right],$
where $\rho_{h}\equiv\lambda_{h}\int_{0}^{\infty}x{\rm d}G(x)\in(0,\infty)$ and we applied: (1) Theorem 1 (see, also the discussion in Section 6.2) in order to bound the ratio $\frac{\mathbb{E}\tau_{\lambda_{h}}}{\mathbb{E}[\tau_{\lambda(\cdot-E)}|E\leq\kappa]}$ from above by one. (2) $\tau_{\lambda(\cdot-E)}$ is not less than the service time of the first customer having distribution $G(\cdot)\equiv G(\Psi)(\cdot)$.
Primarily, by setting $g(w)=w^{k}$ for some $k\in\mathbb{N}$, the proof that the $k$’th moment of the steady-state distribution of $W_{\lambda}(\cdot)$ is finite reduces to showing that the $k$’th moment of the steady-state distribution of $W_{\lambda_{h}}(\cdot)$ is finite.
For future research, it might be interesting to try to evaluate the quality of this bound, i.e., to see whether there are tighter bounds. In addition, it is also interesting to look for lower bounds.
6.4 Tail distribution of $\mathcal{A}_{\lambda}(g)$ and $\mathcal{A}^{*}_{\lambda}(g)$